ADNAN MUHAMMED

Grade 12,

The fourth term of an A.P is equal to 3 times the first term. Its 9th term is 38. Find the 15th term.

The fourth term of an A.P is equal to 3 times the first term. Its 9th term is 38. Find the 15th term.

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
T4 = 3T1 = 3a
T9 = 38 = a + 8d
Also
a + 3d = 3a
3d = 2a
Now
Solve the equations
76 = 2a + 16d
76 = 19d
d = 4
Hence
a = 6
Now 15th term (T15) = a + 14 d = 6 + 56 = 62
 
Regards
Arun(askIITians forum expert)